73.8 is what percent of 90?
Found 2 solutions by macston, stanbon:
Answer by macston(5194)

73.8/90(100%)=82% ANSWER: 73.8 is 82% of 90.

Answer by stanbon(75887)

73.8 is what percent of 90?
-----
x/100 = 73.8/90
x = 7380/90
x = 82
-----
Ans: 82%
-------------
Cheers,
Stan H.
==========